Subject: 3.5-m Users Committee minutes 8/04/08
From: Bruce Gillespie
Submitted: Mon, 18 Aug 2008 16:05:01 -0600
Message number: 1110
(previous: 1109,
next: 1111
up: Index)
APO 3.5-m Users Committee Phonecon, 8/04/08
Attending: Suzanne Hawley, Bruce Gillespie, Mark Klaene, Remy
Indebetouw, Jon Fulbright, John Bally, Michael Strauss, Jon Holtzman,
Scott Anderson, John Wilson, Bill Ketzeback
Absent: Al Harper
Minutes taken by Bruce Gillespie
**********************************
User feedback, comments from institutional representatives:
Remy mentioned that the UVa observers were learning how to use
TripleSpec, which has had a few hiccups. Scott reported on a recent
orientation visit to APO by UW grad students.
**********************************
Discussion of telescope/instruments report - Klaene:
The report is given below. Mark highlighted how the fire danger at
the site has abated due to recent heavy rains. The emergency backup
generator was repaired, as have underground water leaks. During a
small engineering shutdown in July, the top end was inspected and
everything looked good. For the scientific instruments, the DIS was
given a new instrument control computer, and vacuum problems with the
echelle were addressed. Agile will soon be moved to its new permanent
port, with a dedicated rotator. Some on-sky commissioning of
TripleSpec occured, and some minor issues with it are being worked
out. For CIF projects, the new tertiary mirror support system was
assembled at the site; installation will begin shortly. The new axis
controller for the altitude drive is on the docket for the major
summer shutdown, which will start on 6 August and run through 22 August.
Michael Strauss asked about the nature of the work done to the
telescope top end. Mark said that it was routine service and
inspection. Bruce asked what observers might notice after the
shutdown with the new tertiary supports. Mark answered by saying that
the observers should see improvement in telescope pointing stability
and observing efficiency. On auto-focus, Mark added that a fairly
robust version is being tested on the sky, and Bill said that
observers can elect to have auto-focus turned on during some
observations. Its development and testing have been ongoing for about
a month, and more on-sky experience is needed before we put it into
routine use.
*************************
3.5-m Telescope, Instruments, and CIF Projects Highlights, 6/04/08
through 7/30/08
Mark Klaene & Bruce Gillespie
0) Overview
Observing and operations were affected by summer precipitation for
most of this period. Rains fell unexpectedly during June and the
monsoonal flow arrived on schedule (in early July), and was added to
by the remnants of hurricane Dolly. Total rainfall during this period
was nearly 15 inches. The Lincoln National Forest was reopened to the
public on July 2nd.
On site-related matters, all underground water leaks have been
repaired, as well as a fuel leak in the standby generator. Utility
power outages were frequent due to rain and lightning, but the only
storm-related damage has been to the fire alarm system.
SDSS-II observing concluded, and observing started for SDSS-III.
Operational responsibility of 0.5-m telescope was transferred to the
site because the telescope is not needed for SDSS-III.
On August 27th, the Discovery Channel TV "Mythbusters" program is
scheduled to air an episode partially filmed at APO concerning the
APOLLO lunar laser-ranging experiment at APO.
1) Telescope
We started the shutdown activities by servicing the secondary mirror
systems, and moved FastCam to the MB3 mirror-cell port in preparation
for Agile's move to the TR2 mirror-cell port. Updated focus scripts
and new TUI versions were released. A spare computer for the 3.5-m
hub was put into service. Although on-sky time was limited by the
weather, the telescope behaved well during the period.
2) Instruments
DIS, and SPIcam operated smoothly throughout the period. A new spare
DIS instrument control computer was brought on-line.
The echelle had problems with its ion pump and LN2 fill system that
caused inadvertent warm-ups and detector contamination. After several
warming and pumping cycles, the chip contamination is now gone. The
ion pump problems are now resolved.
There has been no recent news from CU regarding the F-P etalon that
was removed from NIC-FIPS and taken back to CU in May. UVa will be
making a newly designed fanout board for NIC-FPS to address noise
issues--work could begin by late September.
TripleSpec commissioning continues with shared-risk use this quarter.
The preliminary user documentation has been updated and reformatted.
Problems with LN2 fills and cart still persist, with some fixes being
worked. Updates to the TUI interface have also been recently
implemented. Procurement of spares and acceptance documentation is
progressing.
Agile has seen a fair amount of use during this period. For Agile's
upgrade and migration to a permanent port, all parts have been
assembled and bench testing is in progress for Agile's move to the TR2
port, with a new rotator. In this permanently mounted configuration,
users will have a choice of 3-inch filters on a SPIcam filter wheel,
or 2-inch filters on a manual slide. The camera will be installed with
a fixed focal reducer.
The visiting instruments (GFP and APOLLO) were both operational. The
IR Camera for APOLLO aircraft-spotting has been moved to the telescope
center boresight.
3) CIF projects
The tertiary support/actuator upgrade is in progress. The new parts
are on the mountain, and they have been assembled and undergone bench
testing that simulates various altitude angles. Work has also begun
on an upgraded motion controller for the altitude axis.
4) Summer Shutdown
On August 6th, we will start a 17-day shutdown for the tertiary mirror
support upgrade, as well as for M1 washing and several additional
maintenance tasks (e.g., annual instrument servicing, etc.).
**********************************
Summer shutdown plans - Klaene
In addition to the report given above, Mark reminded the committee
that when we come out of the summer shutdown on 22 August, weather may
have prevented us from fully re-commissioning the telescope. Although
it looks less likely than in previous years, users should be prepared
that their programs may need to be preempted until enough on-sky time
is available to get the telescope ready for science use.
**********************************
Narrow-band filters for Agile - Hawley:
Suzanne reported that only a couple of users said that they would use
2x2-inch narrow-band filters with Agile. Therefore, the 2-inch filter
slide will be used for these programs, and the action item for this
topic is closed.
**********************************
TripleSpec observing during Q3, and plans for Q4 - Hawley:
Suzanne said that TripleSpec had been block-scheduled in Q3 on a
shared-risk basis; there was one successful run in July, and the
weather interfered with the others. There is one block of observing
scheduled in September. The Q4 scheduling instructions are coming out
soon, and there will be one block in October for TripleSpec, and
thereafter it will be scheduled as a normal instrument. John Wilson
thought that this plan is OK; the shared-risk block observing has been
a good learning experience for the users, and UVa and APO staff
members. The users guide has been updated based on the commissioning
tests. There are minor issues involving background subtraction for
slit viewer images. Also, it has been noted that due to the optical
design, out-of-focus images in the slit viewer show some astigmatism.
**********************************
SDSS-III update - Gillespie, Hawley:
SDSS-III observing began 15 July. More than twenty institutions have
joined the collaboration. The Sloan Foundation has issued its first
payment for operational support, and the newly formed Advisory Council
(AC-III) approved survey spending through 1 September, and will meet
again in August for approval to conduct the survey through the
remainder of the year. The Cooperative Agreement with the NSF is
being negotiated, and the DoE has engaged the NSF and the SDSS-III in
joint discussions on funding support. The MARVELS extra-solar planet
tracker instrument is in final assembly and commissioning at UFla, and
is on track for shipment to APO in early September.
**********************************
New business - Hawley:
Although not on the agenda, Suzanne mentioned that she wanted to
notify the committee that it is time to start collecting Science
Highlights from the past year to be used for our annual report to the
ARC Board of Governors. Please send one or two science highlight from
the telescope users from your institutions to Suzanne (slh at
astro.washington.edu) and Bruce (gillespi at apo.nmsu.edu) not later
than 1 October. We typically try to put the highlights (text,
graphics, and images) on a single slide.
**********************************
ACTION ITEMS:
[open from previous months]:
ACTION: Users committee members and users should contact Suzanne and
Mark Klaene (slh at astro.washington.edu, mark at apo.nmsu.edu) if
they intend to use narrow-band filters for their future Agile science
programs--CLOSED this meeting.
[new actions from this month]:
ACTION: Committee members should send one or two science highlights
from the telescope users from their institutions to Suzanne (slh at
astro.washington.edu) and Bruce (gillespi at apo.nmsu.edu) not later
than 1 October for use in the Annual Report to the ARC Board of
Governors meeting in November.
**********************************
Next meeting:
The next Users Committee phonecon will be on Monday, 15 September, at
8:30 AM Pacific Time. The agenda and other materials will be sent to
the committee members during the preceding week.
APO APO APO APO APO Apache Point Observatory 3.5m APO APO APO
APO
APO This is message 1110 in the apo35-general archive. You can find
APO the archive on http://www.astro.princeton.edu/APO/apo35-general/INDEX.html
APO To join/leave the list, send mail to apo35-request@astro.princeton.edu
APO To post a message, mail it to apo35-general@astro.princeton.edu
APO
APO APO APO APO APO APO APO APO APO APO APO APO APO APO APO APO